OpenStudy (moazzam07):

Trina's Toy Company charges $0.85 per pound to ship toys. Part A: Write an equation to determine the total cost, c, to ship p pounds of toys. Use your equation to determine the cost of shipping 2 pounds of toys. (6 points) Part B: If the company reduces the cost to ship toys by 0.05 per pound, write an equation to determine the total cost, c, to ship p pounds of toys with the reduced cost. (4 points)

OpenStudy (jjwalton04):

Which Part do you not understand?

OpenStudy (moazzam07):

both

OpenStudy (moazzam07):

Can you explain how to do both parts please @ jjwalton04

OpenStudy (moazzam07):

@jjwalton04

OpenStudy (jjwalton04):

ok hold on

OpenStudy (jjwalton04):

Part a=\[0.85p=c\] is the main equation and we need to find the price for two pounds sowe plug in 2 for p and just solve 0.85*2=?

OpenStudy (jjwalton04):

Easy enough right?

OpenStudy (moazzam07):

yes

OpenStudy (jjwalton04):

So when your finished you should get 0.85*2=1.7 Theres A. Now for b, Step one 0.85-0.05 Step two 0.80*2=1.6

OpenStudy (moazzam07):

step one - 0.8 step two -

OpenStudy (moazzam07):

0.80*2=1.6

OpenStudy (jjwalton04):

Yep Thats it.

OpenStudy (moazzam07):

Ok so im done?

OpenStudy (jjwalton04):

Lets see, we explained the price for two pounds and we explained the price for two pounds with the deducted price so...... Yeah

OpenStudy (jjwalton04):

Btw the format is p*x=c

OpenStudy (moazzam07):

thanks

OpenStudy (jjwalton04):

well price*pounds=cost

OpenStudy (moazzam07):

ok thank you

OpenStudy (jjwalton04):

Anytime, if you need me again just tag me